### v1

# Fully heavy tetraquark states with diquark-antidiquark configuration

Abstract
The study of the fully heavy tetraquark state helps to better understand the characteristics of compact exotic hadrons formed by quarks and gluons. In this work, we systematically investigate the mass spectra and decays of fully heavy tetraquark states using a diquark-antidiquark model based on the one gluon exchange process. The tetraquark states are treated as compact diquark–antidiquark configurations, denoted as $QQ][\bar{Q}\bar{Q}]$, where $Q=c$ or $b$. By properly accounting for the Pauli principle, basis wave functions incorporating flavor, color, and spin are constructed. The required model parameters are derived from fitting conventional hadron experimental values and lattice QCD calculations. We calculate the mass spectra and decays of the fully heavy tetraquarks in the lower $S$- and $P$-waves. Our results do not support interpreting $X(6600)$, $X(6900)$, and $X(7200)$ as $S$-wave fully charmed tetraquark states. Nevertheless, our analysis suggests that $X(6600)$ can be a good candidate for the $P$-wave fully charmed tetraquark state. Furthermore, we also identifies several possible fully heavy tetraquark states that may be observable in future experiments.
I Introduction
Quantum Chromodynamics (QCD) allows for a variety of exotic hadronic structures beyond conventional mesons and baryons, such as multiquark states, glueballs, and hybrid states [1, 2, 3]. Among these, tetraquark states have become a major focus in hadron physics. Since the Belle Collaboration’s discovery of the $X(3872)$ resonance in 2003 [4], a growing number of exotic hadrons have been observed experimentally [5]. These discoveries not only confirm the existence of exotic hadronic states but also inspire further theoretical and experimental investigations into their properties. Among the tetraquark states, fully heavy tetraquarks represent a particularly intriguing class. In such systems, light-meson exchange is forbidden, and the interaction is dominated by short-range color forces mediated by gluon exchange, suggesting a compact multiquark configuration.
In recent years, several candidates for fully charmed tetraquark states have been observed experimentally. In 2020, the LHCb Collaboration observed a broad structure above twice the $J/\psi$ mass threshold ranging from 6.2 to 6.8 GeV and a narrower structure at about 6.9 GeV, called $X(6900)$ [6]. Very recently, the CMS and ATLAS Collaborations also researched the di-$J/\psi$ channel and both of them confirmed the existence of the $X(6900)$ [7, 8]. Besides, the CMS Collaboration reported two new structures, the $X(6600)$ and $X(7200)$, in the di-$J/\psi$ invariant mass spectrum [7]. The ATLAS Collaboration also observed the new structures, $X(6400)$ and $X(6600)$, in the di-$J/\psi$ channel and $X(7200)$ in the $J/\psi$-$\psi(2S)$ invariant mass spectrum [8]. These experiments have motivated a large number of theoretical studies on fully charmed tetraquark states [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Many theoretical studies suggest that $X(6900)$ is a $2S$-wave tetraquark state [23, 24, 25, 26, 27, 28, 16], while others propose it as an $S$-wave state with angular quantum number $L=0$ and total quark spin $S=2$ [29]. Some works interpret $X(6600)$ as a $2S$-wave state [28, 30], whereas others regard it as a $1S$-wave [23]. There are also studies that suggest $X(7200)$ could be a $3S$-wave [27, 28] or a $2P$-wave state [16].
Furthermore, some new experimental evidence for fully bottom tetraquark states has been reported. In 2016, the CMS Collaboration announced the first observation of $\Upsilon(1S)$ pair production in $pp$ collisions at $\sqrt{s}=8$ TeV [31]. They found an excess with a global significance of 3.6 $\sigma$ in the $4\mu$ channel. The invariant mass of this excess is $18.4\pm 0.1\pm 0.2$ GeV [32]. Two years later, the LHCb Collaboration performed a research for the decay of the beautiful tetraquark, $X$, to the $\Upsilon(1S)\mu^{+}\mu^{-}$ final state. However, no significant excess is seen for any mass hypothesis in the range [17.5, 20.0][17.5,\ 20.0] GeV [33]. In 2019, other production of double $\Upsilon(1S)$ is observed in Cu + Au collisions at the Relativistic Heavy Ion Collider by ANDY Collaboration. They shows evidence of a significant peak at $M=18.12\pm 0.15\pm 0.6$ GeV [34], which is compatible with the first observation of an all-$b$ tetraquark. The fully bottom tetraquark states have also been the subject of extensive theoretical investigations [9, 35, 36, 37, 20]. Given that only a few experiments have observed these possible fully bottom tetraquark states, further experimental investigations are required to confirm their existence.
Fully heavy tetraquark states, especially the fully charmed and fully bottom tetraquark states, have attracted extensive attention from theorists. A wide range of theoretical approaches have been employed to study them, including the QCD sum rules [38, 10, 39, 11], a variety of quark models [24, 13, 16, 40, 18, 20, 41, 42, 9, 43], the diquark models [12, 35, 19, 25], the chromomagnetic model [44, 45, 46], the diffusion Monte Carlo approach [36, 47, 48], and the other phenomenological models [26, 27, 28, 29, 30]. A few studies suggest that the fully heavy tetraquark states $cc\bar{c}\bar{c}$ and $bb\bar{b}\bar{b}$ could be bound states. In Ref. [46], the lowest mass of the $cc\bar{c}\bar{c}$ system lies below the di-$\eta_{c}$ threshold, while the lightest S-wave $bb\bar{b}\bar{b}$ state appears below the $\eta_{b}\eta_{b}$ threshold in Ref. [35, 49, 9]. However, the majority of investigations conclude that the $S$-wave fully heavy tetraquarks are located above the corresponding double-meson thresholds. For example, in Ref. [50], the authors used the first-principles lattice nonrelativistic QCD methodology to study the low-lying spectrum of the $bb\bar{b}\bar{b}$ tetraquark states with quantum numbers $0^{++}$, $1^{+-}$ and $2^{++}$. However, they found no evidence of a QCD bound tetraquark below the lowest non-interacting thresholds in the channels studied. It is evident that theorists hold diverse views regarding the new experimental observations, offering a variety of interpretations and even arriving at some contradictory conclusions. At present, there is no theoretical explanation that has gained universal acceptance. The search for new theoretical approaches remains an active and ongoing endeavor.
Nowadays, a variety of new exotic hadrons are continuously being observed in experiments, while theoretical progress remains relatively limited. Further theoretical studies are still required to deepen our understanding of the mechanisms underlying strong interactions. In this work, we investigate the mass spectra of fully heavy tetraquark states within the diquark–antidiquark configuration and analyze their strong decay behaviors.
This paper is organized as follows. In Sec. II, we present the effective Hamiltonian employed in this work and provide the basis wave functions corresponding to different quantum numbers. Subsequently, we introduce the formulae used for calculating the strong decay widths. In Sec. III, we obtain the model parameters required for the calculations and describe the fitting methods used to obtain these parameters. We then show our calculated results and provide an analysis of the decay processes. We summarize in Sec. IV. The nature units $c=\hbar=1$ are used throughout.
II Theoretical model
II.1 Effective Hamiltonian
The bound state system formed by four fully heavy quarks can be seen as a unique example of direct interaction between quarks and gluons. Consequently, the interaction in this system is mainly manifested as the one gluon exchange process [51, 52]. For the fully heavy tetraquark systems, the improved nonrelativistic effective Hamiltonian reads [53, 54]
$H=\sum_{i=1}^{4}m_{i}+H_{SS}+H_{SL}+H_{L},$ (1)
where
$H_{SS}={}$ $\sum_{i<j}2\kappa_{ij}(\bm{S}_{i}\cdot\bm{S}_{j}),$ (2)
$H_{SL}={}$ $2A\left(\bm{S}_{\mathcal{Q}}\cdot\bm{L}+\bm{S}_{\bar{\mathcal{Q}}^{\prime}}\cdot\bm{L}\right),$
$H_{L}={}$ $B\frac{L(L+1)}{2}.$
Here, $m_{i}$ denotes the mass of the $i$-th constituent quark mass. $H_{SS}$ represents the interaction of each quark pair in the tetraquark, $H_{SL}$ and $H_{L}$ describe spin-orbit coupling and pure orbital contribution, respectively. $\bm{S}_{\mathcal{Q}}$ in $H_{SL}$ represents the spin operator of the diquark configuration $\mathcal{Q}=[Q_{1}Q_{2}]$. The orbital angular momentum $\bm{L}$ describes the relative motion between the diquark $\mathcal{Q}$ and the anti-diquark $\bar{\mathcal{Q}}^{\prime}$. The coefficients $A$ and $B$ are phenomenological parameters to be determined from experimental data.
II.2 Basis wave functions
To obtain the mass spectra of the fully heavy tetraquark states, we need to determine their wave functions. Here we only show the wave functions of ground states and we represent them as the form $|{(Q_{1}Q_{2})_{s}^{c}(\bar{Q}_{3}\bar{Q}_{4})_{s}^{c}}\rangle$, where the superscript denotes the color and the subscript denotes the spin.
Considering the Pauli principle, if the two quarks in a diquark have the same flavor, the spin and color of the diquark must have opposite symmetries. Therefore, we define an anti-Kronecker delta
$\bar{\delta}_{ij}=\left\{\begin{matrix}0\quad\text{if}\ i=j,\\ 1\quad\text{if}\ i\neq j.\end{matrix}\right$. (3)
The wave functions of the ground states are as follows, (1) $J^{P}=0^{+}$
$\alpha_{1}=$ $|(Q_{1}Q_{2})_{0}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{3}\rangle\bar{\delta}_{12}\bar{\delta}_{34},$ (4)
$\alpha_{2}=$ $|(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{\bar{6}}\rangle,$
$\alpha_{3}=$ $|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle,$
$\alpha_{4}=$ $|(Q_{1}Q_{2})_{1}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{\bar{6}}\rangle\bar{\delta}_{12}\bar{\delta}_{34}.$
(2) $J^{P}=1^{+}$
$\beta_{1}=$ $|(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{\bar{6}}\rangle\bar{\delta}_{34},$ (5)
$\beta_{2}=$ $|(Q_{1}Q_{2})_{0}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle\bar{\delta}_{12},$
$\beta_{3}=$ $|(Q_{1}Q_{2})_{1}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{\bar{6}}\rangle\bar{\delta}_{12},$
$\beta_{4}=$ $|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{3}\rangle\bar{\delta}_{34},$
$\beta_{5}=$ $|(Q_{1}Q_{2})_{1}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{\bar{6}}\rangle\bar{\delta}_{12}\bar{\delta}_{34},$
$\beta_{6}=$ $|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle.$
(3) $J^{P}=2^{+}$
$\gamma_{1}=$ $|(Q_{1}Q_{2})_{1}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{\bar{6}}\rangle\bar{\delta}_{12}\bar{\delta}_{34},$ (6)
$\gamma_{2}=$ $|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle.$
The $\bar{\delta}_{12}$ ($\bar{\delta}_{34}$) means that if $Q_{1}Q_{2}$ ($\bar{Q}_{3}\bar{Q}_{4}$) have the same flavor then this state can not exist. We also consider the charge conjugation and excited states with $L=1$. The complete wave functions are listed in Appendix A.
II.3 Decay width
By inserting the basis vectors of the wave function into the Hamiltonian, one can obtain the mass spectra of the tetraquark states. We can then calculate their decays into different meson final states.
For a particle of mass $M$ and four-momentum $P$ decaying into two particles with momenta $p_{1}$ and $p_{2}$, the partial decay rate is written as
$d\Gamma=\frac{(2\pi)^{4}}{2M}|\mathcal{A}|^{2}d\Phi(P;p_{1},p_{2}),$ (7)
where $\mathcal{A}$ is the amplitude connecting the initial state to the final state and $d\Phi$ is an element of 2-body phase space given by
$d\Phi(P;p_{1},p_{2})$ $=\delta^{4}(P-p_{1}-p_{2})$ (8)
$\times\frac{d^{3}p_{1}}{(2\pi)^{3}2E_{1}}\frac{d^{3}p_{2}}{(2\pi)^{3}2E_{2}}.$
In the Center of Momentum Frame, considering four-momentum conservation in the decay process, we have
$|\bm{p}_{1}|=|\bm{p}_{2}|=\frac{1}{2M}\sqrt{\lambda(M^{2},m_{1}^{2},m_{2}^{2})}$ (9)
where $\lambda(\alpha,\beta,\gamma)=\alpha^{2}+\beta^{2}+\gamma^{2}-2\alpha\beta-2\alpha\gamma-2\beta\gamma$ is the Källén function. Thus, the partial decay rate can be written as
$d\Gamma=\frac{|\mathcal{A}|^{2}}{32\pi^{2}}\frac{|\bm{p}_{1}|}{M^{2}}d\Omega.$ (10)
Then, by integrating the differentiation above, we obtain
$\Gamma=\frac{|\mathcal{A}|^{2}|\bm{p}_{1}|}{8\pi M^{2}}.$ (11)
III results and discussion
III.1 Model parameters
To compute the ground-state masses of fully heavy tetraquark states, we need to determine the constituent masses of the charm and bottom quarks, as well as the parameters $\kappa_{ij}$ associated with quark pairs. We obtain the constituent masses and the parameters $\kappa_{ij}$ from known conventional hadrons. For conventional hadrons that have not yet been observed experimentally, we use data from Lattice QCD [55, 56].
The color states associated with the interaction parameters $\kappa_{ij}$ obtained through conventional hadron spectrum must exclusively correspond to either 1 or $\bar{3}$. Thus we first determine $(\kappa_{Q\bar{Q}})_{1}$ with mesons and $(\kappa_{QQ})_{\bar{3}}$ with baryons.
For mesons with $L=0$, Eq. (1) provides the foundation for deriving
$M=m_{Q_{1}}+m_{Q_{2}}+(\kappa_{Q_{1}\bar{Q}_{2}})_{1}[J(J+1)-\frac{3}{2}].$ (12)
With these equation we can calculate the value of $(\kappa_{Q_{1}\bar{Q}_{2}})_{1}$. Their numerical results are tabulated in Table 1.
Table 1: Parameter $\kappa$ (in MeV) for $q\bar{q}$ with color configuration of singlet and for qqqq with color anti-triplet. And these parameters are derived from fitting the $S$-wave mesons and baryons.
$(\kappa_{c\bar{c}})_{1}$ $(\kappa_{b\bar{b}})_{1}$ $(\kappa_{b\bar{c}})_{1}$ $(\kappa_{cc})_{\bar{3}}$ $(\kappa_{bb})_{\bar{3}}$ $(\kappa_{bc})_{\bar{3}}$
56 31 32 7 4 11
To determine $(\kappa_{QQ})_{\bar{3}}$ with baryons, Eq. (1) can be reformulated as
$M=$ $\sum_{i=1}^{3}m_{i}+2\kappa_{q_{1}q_{2}}(\bm{s}_{q_{1}}\cdot\bm{s}_{q_{2}})+\kappa_{q_{1}q_{3}}(\bm{s}_{q_{1}}\cdot\bm{s}_{q_{3}})$
$+$ $\kappa_{q_{2}q_{3}}(\bm{s}_{q_{2}}\cdot\bm{s}_{q_{3}}).$ (13)
For simplicity, we choose baryons that contain two quarks of the same flavor for our calculation. In this case, Eq. (13) can be written as
$M=$ $2m_{q_{1}}+m_{q_{2}}+(\kappa_{q_{1}q_{1}})_{\bar{3}}[S(S+1)-\frac{3}{2}]$
$+$ $(\kappa_{q_{1}q_{2}})_{\bar{3}}[J(J+1)-S(S+1)-\frac{3}{4}],$ (14)
where $J$ is the $S$ represents the spin of two same-flavor quarks within a baryon.
By considering baryons composed of nnqnnq with quantum numbers $(S,J)$ equal to $(0,\frac{1}{2})$, $(1,\frac{1}{2})$, and $(1,\frac{3}{2})$, the values of ${(\kappa_{nn})}_{\bar{3}}$ and ${(\kappa_{nq})}_{\bar{3}}$ can be determined. Here $n=u,d$ and $q=s,c,b$. We can get a proportional relationship read ${(\kappa_{nn})}_{\bar{3}}/{(\kappa_{nq})}_{\bar{3}}=m_{q}/m_{n}$. Furthermore, the values of ${(\kappa_{bb})}_{\bar{3}}$, ${(\kappa_{cb})}_{\bar{3}}$ can be extracted from $\Omega_{bbb}$, $\Omega_{bbc}$ and $\Omega_{bbc}^{*}$. We can extract ${(\kappa_{cc}})_{\bar{3}}$ from $\Omega_{ccc}$, $\Omega_{ccs}$ and $\Omega_{ccs}^{*}$. All the coefficients are shown in Table 1.
At this juncture, it remains crucial to determine the value of $(\kappa_{QQ})_{6}$, along with establishing the corresponding $(\kappa_{Q\bar{Q}})_{T}$ parameter within the tetraquark configuration. From one-gluon-exchange, we obtain
$(\kappa_{QQ})_{6}={}$ $-\frac{1}{2}(\kappa_{QQ})_{\bar{3}},$ (15)
$(\kappa_{Q\bar{Q}})_{\bar{3}}={}$ $\frac{1}{4}(\kappa_{Q\bar{Q}})_{1},$
$(\kappa_{Q\bar{Q}})_{6}={}$ $\frac{5}{8}(\kappa_{Q\bar{Q}})_{1}.$
where $(\kappa_{Q\bar{Q}})_{\bar{3}}$ means the $\kappa_{Q\bar{Q}}$ for quark-antiquark in tetraquark $|(Q_{1}Q_{2})^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})^{3}\rangle$ and $(\kappa_{Q\bar{Q}})_{6}$ for $|(Q_{1}Q_{2})^{6}(\bar{Q}_{3}\bar{Q}_{4})^{\bar{6}}\rangle$.
The masses of charm and bottom quark we derived are $m_{c}=1556$MeV and $m_{b}=4755$MeV.
After computing the ground state masses, we can include the contributions from $H_{SL}$ and $H_{L}$ to calculate the masses of the $L=1$ excited states. By applying Eq. (1) to charmonium mesons with $S=1$ and $L=0,1$, specifically $J/\psi$, $\chi_{c0}$, $\chi_{c1}$, and $\chi_{c2}$, we can fit the coefficients in $H_{SL}$ and $H_{L}$, we get $A_{cc}=21.83$ MeV and $B_{cc}=426.06$ MeV. Similarly, using the bottomonium mesons $\Upsilon$, $\chi_{b0}$, $\chi_{b1}$, and $\chi_{b2}$, we obtain $A_{bb}=8.23$ MeV and $B_{bb}=438.72$ MeV. For tetraquark states containing both charm and bottom quarks, we use the average values of $A_{cc}$ and $A_{bb}$ (as well as $B_{cc}$ and $B_{bb}$) for our calculations.
Since the tetraquark states discussed in this work are structured as diquark–antidiquark systems, quark rearrangement is required for them to strongly decay into two mesons. The decay widths are calculated using Eq. (11) with $\mathcal{A}$ = 2600 MeV. The value of $\mathcal{A}$ can be obtained by fitting light scalar meson decays [57]. As higher-order decays (such as $P$-wave and $D$-wave decays) contribute less significantly to the total decay width, our analysis focuses primarily on $S$-wave decays. However, we also analyze and discuss $P$-wave decays in special cases.
(a) $cc][\bar{c}\bar{c}]$
(b) $bb][\bar{b}\bar{b}]$
(c) $cb][\bar{c}\bar{b}]$
Figure 1: The mass spectra of the fully-heavy tetraquark states: (a) $cc][\bar{c}\bar{c}]$, (b) $bb][\bar{b}\bar{b}]$, and (c) $cb][\bar{c}\bar{b}]$, each with different quantum numbers. The dash-dotted lines represent the corresponding meson-meson thresholds. In Fig. (c), the “several channels” include many threshold lines; for clarity and visual neatness, those lines overlapping with the energy levels have been omitted.
(a) $cc][\bar{c}\bar{b}]$
(b) $bb][\bar{b}\bar{c}]$
(c) $cc][\bar{b}\bar{b}]$
Figure 2: The mass spectra of the fully-heavy tetraquark states: (a) $cc][\bar{c}\bar{b}]$, (b) $bb][\bar{b}\bar{c}]$, and (c) $cc][\bar{b}\bar{b}]$, each with different quantum numbers.The dash-dotted lines represent the corresponding meson-meson thresholds.
Table 2: The quantum numbers, masses, eigenvectors of the basis wave functions, decay channels of $cc][\bar{c}\bar{c}]$ and $bb][\bar{b}\bar{b}]$ configurations
system $J^{PC}$ mass eigenvector Decay channels $\Gamma$(MeV)]
$cc][\bar{c}\bar{c}]$ $0^{++}$ 6098.45 (0.60,0.80) $\eta_{c}\eta_{c}[4.53]$
6314.21 (0.80,0.60)(-0.80,0.60) $\eta_{c}\eta_{c}[6.95];J/\psi J/\psi[4.14]$
$1^{+-}$ 6204.49 1 $\eta_{c}J/\psi[4.31]$
$2^{++}$ 6260.89 1 $J/\psi J/\psi[3.14]$
$0^{-+}$ 6543.23 1 $\eta_{c}\chi_{c0}[4.29]$
$1^{--}$ 6524.51 (0.60,0.80) $\eta_{c}h_{c}[1.39];J/\psi\chi_{c0}[1.29]$
6740.27 (0.80,0.60)(-0.80,0.60) $\eta_{c}h_{c}[5.16];J/\psi\chi_{c0}[5.15];J/\psi\chi_{c1}[3.93];J/\psi\chi_{c2}[3.19]$
6555.97 1 $\eta_{c}h_{c}[2.43];J/\psi\chi_{c0}[2.38]$
$1^{-+}$ 6586.89 1 $\eta_{c}\chi_{c1}[3.39]$
$bb][\bar{b}\bar{b}]$ $0^{++}$ 18948.9 (0.601,0.799) $\eta_{b}\eta_{b}[0.90];\Upsilon\Upsilon[0.39]$
19066.9 (0.799,0.601)(-0.799,0.601) $\eta_{b}\eta_{b}[1.18];\Upsilon\Upsilon[0.87]$
$1^{+-}$ 19006.9 1 $\eta_{b}\Upsilon[0.88]$
$2^{++}$ 19037.7 1 $\Upsilon\Upsilon[0.78]$
$0^{-+}$ 19412.7 1 $\eta_{b}\chi_{b0}[0.87];\Upsilon h_{b}[0.51]$
$1^{--}$ 19387.6 (0.601,0.799) $\eta_{b}h_{b}[0.67];\Upsilon\chi_{b0}[0.58];\Upsilon\chi_{b1}[0.41];\Upsilon\chi_{b2}[0.27]$
19505.6 (0.799,0.601)(-0.799,0.601) $\eta_{b}h_{b}[1.00];\Upsilon\chi_{b0}[0.95];\Upsilon\chi_{b1}[0.86];\Upsilon\chi_{b2}[0.80]$
19427.0 1 $\eta_{b}h_{b}[0.80];\Upsilon\chi_{b0}[0.73];\Upsilon\chi_{b1}[0.60];\Upsilon\chi_{b2}[0.52]$
$1^{-+}$ 19429.2 1 $\eta_{b}$$\chi_{b1}[0.82];\Upsilon h_{b}[0.58]$
Table 3: The quantum numbers, masses, eigenvectors of the basis wave functions and decay channels of $cb][\bar{c}\bar{b}]$ configuration. For certain states, there are no available channels that both satisfy the conservation laws and have thresholds below the mass of the state.
$J^{PC}$ MM(MeV) eigenvector Decay channels $\Gamma$(MeV)]
$0^{++}$ 12477.8 (0.15,0.52,0.82,0.20)(-0.15,-0.52,-0.82,-0.20) $\eta_{c}\eta_{b}[1.14]$
12548.6 (0.57,0.20,0.17,0.78)(0.57,-0.20,-0.17,0.78) $\eta_{c}\eta_{b}[1.49]$
12690.1 (0.81,0.06,0.04,0.59)(-0.81,-0.06,0.04,0.59) $\eta_{c}\eta_{b}[2.00];J/\psi\Upsilon[1.32];B_{c}\bar{B}_{c}[1.58];B_{c}^{*}\bar{B}_{c}^{*}[0.70]$
12633.4 (0.02,0.83,0.55,0.10)(-0.02,0.83,-0.55,0.10) $\eta_{c}\eta_{b}[1.82];J/\psi\Upsilon[1.01];B_{c}\bar{B}_{c}[1.23]$
$1^{++}$ 12665.6 (0.79,0.61)(-0.79,-0.61) $J/\psi\Upsilon[1.20];B_{c}\bar{B}_{c}^{*}[1.03];B_{c}^{*}\bar{B}_{c}^{*}[0.25]$
12582.8 (0.61,0.79)(0.61,-0.79) $J/\psi\Upsilon[0.59]$
$1^{+-}$ 12547.4 (0.46,0.48,0.71,0.21)(0.46,0.48,-0.71,-0.21) $\eta_{c}\Upsilon[1.17];J/\psi\eta_{b}[0.84]$
12655.4 (0.77,0.62,0.04,0.15)(0.77,-0.62,0.04,0.15) $\eta_{c}\Upsilon[1.66];J/\psi\eta_{b}[1.46];B_{c}\bar{B}_{c}^{*}[0.94]$
12590.8 (0.43,0.53,0.70,0.20)(0.43,0.53,0.70,-0.20) $\eta_{c}\Upsilon[1.39];J/\psi\eta_{b}[1.13]$
12617.6 (0.07,0.32,0.01,0.94)(0.07,0.32,-0.01,0.94) $\eta_{c}\Upsilon[1.51];J/\psi\eta_{b}[1.28];B_{c}\bar{B}_{c}^{*}[0.47]$
$2^{++}$ 12666.1 (0.91,0.41)(-0.91,-0.41) $J/\psi\Upsilon[1.20];B_{c}^{*}\bar{B}_{c}^{*}[0.27]$
12649.1 (0.41,0.91)(0.41,-0.91) $J/\psi\Upsilon[1.11]$
$0^{--}$ 13037.9 (0.79,0.61)(-0.79,-0.61) $J/\psi\chi_{b1}[0.76];\chi_{c1}\Upsilon[0.93];B_{c}\bar{B}_{c0}[0.91]$
12955.1 (0.61,0.79)(0.61,-0.79)
$0^{-+}$ 12919.7 (0.46,0.48,0.71,0.21)(0.46,0.48,-0.71,-0.21) $\eta_{c}\chi_{b0}[0.96];\chi_{c0}\eta_{b}[1.18]$
13027.7 (0.77,0.62,0.04,0.15)(0.77,-0.62,0.04,0.15) $\eta_{c}\chi_{b0}[1.47];\chi_{c0}\eta_{b}[1.66];J/\psi h_{b}[0.61];h_{c}\Upsilon[0.74];B_{c}\bar{B}_{c0}[0.82]$
12963.1 (0.43,0.53,0.70,0.20)(0.43,0.53,0.70,-0.20) $\eta_{c}\chi_{b0}[1.19];\chi_{c0}\eta_{b}[1.39]$
12989.9 (0.07,0.32,0.01,0.94)(0.07,0.32,-0.01,0.94) $\eta_{c}\chi_{b0}[1.31];\chi_{c0}\eta_{b}[1.51]$
$1^{--}$ 12910.2 (0.15,0.52,0.82,0.20)(-0.15,-0.52,-0.82,-0.20) $\eta_{c}\eta_{b}[0.57];\chi_{c0}\Upsilon[0.68]$
12981 (0.57,0.20,0.17,0.78)(0.57,-0.20,-0.17,0.78) $\eta_{c}\eta_{b}[1.66];J/\psi\chi_{b0}[1.39];\chi_{c0}\Upsilon[1.76];\chi_{c1}\Upsilon[1.38];h_{c}\eta_{b}[1.59]$
13122.5 (0.81,0.06,0.04,0.59)(-0.81,-0.06,0.04,0.59) $\eta_{c}\eta_{b}[1.66];J/\psi\chi_{b0}[1.39];J/\psi\chi_{b1}[1.21];J/\psi\chi_{b2}[1.15];\chi_{c0}\Upsilon[1.76];\chi_{c1}\Upsilon[1.38]$
$\chi_{c2}\Upsilon[1.16];h_{c}\eta_{b}[1.59];B_{c}\bar{B}_{c1}[1.34];B_{c}^{*}\bar{B}_{c0}[1.13];B_{c}^{*}\bar{B}_{c1}[0.94]$
13065.8 (0.02,0.83,0.55,0.10)(-0.02,0.83,-0.55,0.10) $\eta_{c}\eta_{b}[1.45];J/\psi\chi_{b0}[1.14];J/\psi\chi_{b1}[0.91];J/\psi\chi_{b2}[0.82];\chi_{c0}\Upsilon[1.55];\chi_{c1}\Upsilon[1.10]$
$\chi_{c2}\Upsilon[0.80];h_{c}\eta_{b}[1.35];B_{c}\bar{B}_{c1}[0.95];B_{c}^{*}\bar{B}_{c0}[0.61]$
13067.9 (0.79,0.61)(-0.79,-0.61) $\eta_{c}\eta_{b}[1.46];J/\psi\chi_{b0}[1.15];J/\psi\chi_{b1}[0.92];J/\psi\chi_{b2}[0.83];\chi_{c0}\Upsilon[1.56];\chi_{c1}\Upsilon[1.11]$
$\chi_{c2}\Upsilon[0.81];h_{c}\eta_{b}[1.36];B_{c}\bar{B}_{c1}[0.96];B_{c}^{*}\bar{B}_{c0}[0.63];B_{c}^{*}\bar{B}_{c1}[0.12]$
12985.1 (0.61,0.79)(0.61,-0.79) $\eta_{c}\eta_{b}[1.09];J/\psi\chi_{b0}[0.59];\chi_{c0}\Upsilon[1.19];\chi_{c1}\Upsilon[0.43];h_{c}\eta_{b}[0.89]$
13008.3 (0.91,0.41)(-0.91,-0.41) $\eta_{c}\eta_{b}[1.21];J/\psi\chi_{b0}[0.79];J/\psi\chi_{b1}[0.38];\chi_{c0}\Upsilon[1.31];\chi_{c1}\Upsilon[0.70];h_{c}\eta_{b}[1.05]$
12991.4 (0.41,0.91)(0.41,-0.91) $\eta_{c}\eta_{b}[1.13];J/\psi\chi_{b0}[0.65];\chi_{c0}\Upsilon[1.22];\chi_{c1}\Upsilon[0.51];h_{c}\eta_{b}[0.94]$
$1^{-+}$ 12949.7 (0.46,0.48,0.71,0.21)(0.46,0.48,-0.71,-0.21) $\eta_{c}\chi_{b1}[0.93];\chi_{c1}\eta_{b}[0.73]$
13057.7 (0.77,0.62,0.04,0.15)(0.77,-0.62,0.04,0.15) $\eta_{c}\chi_{b1}[1.45];J/\psi h_{b}[0.85];\chi_{c1}\eta_{b}[1.38];h_{c}\Upsilon[0.96];B_{c}\bar{B}_{c1}[0.87];B_{c}^{*}\bar{B}_{c0}[0.49]$
12993.1 (0.43,0.53,0.70,0.20)(0.43,0.53,0.70,-0.20) $\eta_{c}\chi_{b1}[1.17];\chi_{c1}\eta_{b}[1.05];h_{c}\Upsilon[0.31]$
13019.9 (0.07,0.32,0.01,0.94)(0.07,0.32,-0.01,0.94) $\eta_{c}\chi_{b1}[1.29];J/\psi h_{b}[0.53];\chi_{c1}\eta_{b}[1.20];h_{c}\Upsilon[0.67];B_{c}\bar{B}_{c1}[0.39]$
Table 4: The quantum numbers, masses, eigenvectors of the basis wave functions and decay channels of $cc][\bar{c}\bar{b}]$ configuration
$J^{P}$ MM(MeV) eigenvector Decay channels $\Gamma$(MeV)]
$0^{+}$ 9328.67 (0.597,0.803) $\eta_{c}B_{c}[1.65]$
9497.82 (0.803,0.597)(-0.803,0.597) $\eta_{c}B_{c}[2.96];J/\psi B_{c}^{*}[1.61]$
$1^{+}$ 9363.47 (0.622,0.690,0.369)(0.622,-0.690,0.369) $J/\psi B_{c}[1.36]$
9467.10 (0.774,0.613,0.160)(0.774,0.613,-0.160) $\eta_{c}B_{c}^{*}[1.90];J/\psi B_{c}[2.37];J/\psi B_{c}^{*}[1.21]$
9416.91 (0.116,0.385,0.916)(-0.116,0.385,0.916) $\eta_{c}B_{c}^{*}[1.32];J/\psi B_{c}[1.96]$
$2^{+}$ 9454.99 1 $J/\psi B_{c}^{*}[1.01]$
$0^{-}$ 9735.74 (0.622,0.690,0.369)(0.622,-0.690,0.369) $\eta_{c}B_{c0}[0.29];J/\psi B_{c1}[0.15]$
9839.37 (0.774,0.613,0.160)(0.774,0.613,-0.160) $\eta_{c}B_{c0}[2.28];J/\psi B_{c1}[2.15];\chi_{c0}B_{c}[0.46]$
9789.18 (0.116,0.385,0.916)(-0.116,0.385,0.916) $\eta_{c}B_{c0}[1.87];J/\psi B_{c1}[1.75]$
$1^{-}$ 9761.06 (0.597,0.803) $J/\psi B_{c1}[0.74];h_{c}B_{c}^{*}[1.16]$
9930.21 (0.803,0.597)(-0.803,0.597) $\eta_{c}B_{c1}[2.22];J/\psi B_{c0}[2.10];J/\psi B_{c1}[2.48];\chi_{c0}B_{c}^{*}[1.73];\chi_{c1}B_{c}[1.58]$
$\chi_{c1}B_{c}^{*}[1.21];h_{c}B_{c}[1.97];h_{c}B_{c}^{*}[2.57];\chi_{c2}B_{c}^{*}[1.76]$
9765.8 (0.622,0.690,0.369)(0.622,-0.690,0.369) $J/\psi B_{c1}[0.84];h_{c}B_{c}^{*}[1.23]$
9869.43 (0.774,0.613,0.160)(0.774,0.613,-0.160) $\eta_{c}B_{c1}[1.71];J/\psi B_{c0}[1.55];J/\psi B_{c1}[2.05];\chi_{c0}B_{c}^{*}[0.98];\chi_{c1}B_{c}[0.67]$
$h_{c}B_{c}[1.40];h_{c}B_{c}^{*}[2.18];\chi_{c2}B_{c}^{*}[1.09]$
9819.24 (0.116,0.385,0.916)(-0.116,0.385,0.916) $\eta_{c}B_{c1}[1.09];J/\psi B_{c0}[0.83];J/\psi B_{c1}[1.60];h_{c}B_{c}[0.58];h_{c}B_{c}^{*}[1.79]$
9797.2 1 $\eta_{c}B_{c1}[0.65];J/\psi B_{c1}[1.34];h_{c}B_{c}^{*}[1.59]$
Table 5: The quantum numbers, masses, eigenvectors of the basis wave functions and decay channels of $bb][\bar{b}\bar{c}]$ configuration
$J^{P}$ MM(MeV) eigenvector Decay channels $\Gamma$(MeV)]
$0^{+}$ 15753.9 $(0.595,0.803)$ $\eta_{b}\bar{B}_{c}[0.85]$
15874.1 (0.803,0.595)(-0.803,0.595) $\eta_{b}\bar{B}_{c}[1.32];\Upsilon\bar{B}_{c}*[0.85]$
$1^{+}$ 15778.6 (0.624,0.782,0.02)(0.624,-0.782,-0.02) $\eta_{b}\bar{B}_{c}^{*}[0.66];\Upsilon\bar{B}_{c}[0.62]$
15846.7 $(0.782,0.624,0.010)$ $\eta_{b}\bar{B}_{c}^{*}[1.01];\Upsilon\bar{B}_{c}[0.99];\Upsilon\bar{B}_{c}^{*}[0.69]$
15812.1 (0.005,0.022,0.999)(0.005,-0.022,0.999) $\eta_{b}\bar{B}_{c}^{*}[0.85];\Upsilon\bar{B}_{c}[0.82];\Upsilon\bar{B}_{c}^{*}[0.43]$
$2^{+}$ 15843.4 1 $\Upsilon\bar{B}_{c}^{*}[0.67]$
$0^{-}$ 16150.9 (0.624,0.782,0.02)(0.624,-0.782,-0.02) $\eta_{b}\bar{B}_{c0}[0.58];\chi_{b0}\bar{B}_{c}[0.37]$
16219 $(0.782,0.624,0.010)$ $\eta_{b}\bar{B}_{c0}[0.94];\Upsilon\bar{B}_{c1}[0.43];\chi_{b0}\bar{B}_{c}[0.83]$
16184.4 (0.005,0.022,0.999)(0.005,-0.022,0.999) $\eta_{b}\bar{B}_{c0}[0.78];\chi_{b0}\bar{B}_{c}[0.64]$
$1^{-}$ 16186.3 $(0.595,0.803)$ $\eta_{b}\bar{B}_{c1}[0.65];\Upsilon\bar{B}_{c0}[0.34];\chi_{b1}\bar{B}_{c}[0.39];h_{b}\bar{B}_{c}[0.32]$
16306.5 (0.803,0.595)(-0.803,0.595) $\eta_{b}\bar{B}_{c1}[1.18];\Upsilon\bar{B}_{c0}[1.04];\Upsilon\bar{B}_{c1}[0.94];\chi_{b0}\bar{B}_{c}^{*}[0.96];\chi_{b1}\bar{B}_{c}[1.05]$
$\chi_{b1}\bar{B}_{c}^{*}[0.81];h_{b}\bar{B}_{c}[1.02];h_{b}\bar{B}_{c}^{*}[0.78];\chi_{b2}\bar{B}_{c}^{*}[0.71]$
16180.9 (0.624,0.782,0.02)(0.624,-0.782,-0.02) $\eta_{b}\bar{B}_{c1}[0.62];\Upsilon\bar{B}_{c0}[0.27];\chi_{b1}\bar{B}_{c}[0.33];h_{b}\bar{B}_{c}[0.24]$
16249 $(0.782,0.624,0.010)$ $\eta_{b}\bar{B}_{c1}[0.97];\Upsilon\bar{B}_{c0}[0.79];\Upsilon\bar{B}_{c1}[0.66];\chi_{b0}\bar{B}_{c}^{*}[0.69];\chi_{b1}\bar{B}_{c}[0.81]$
$\chi_{b1}\bar{B}_{c}^{*}[0.45];h_{b}\bar{B}_{c}[0.78];h_{b}\bar{B}_{c}^{*}[0.39];\chi_{b2}\bar{B}_{c}^{*}[0.22]$
16214.4 (0.005,0.022,0.999)(0.005,-0.022,0.999) $\eta_{b}\bar{B}_{c1}[0.81];\Upsilon\bar{B}_{c0}[0.59];\Upsilon\bar{B}_{c1}[0.39];\chi_{b0}\bar{B}_{c}^{*}[0.44];\chi_{b1}\bar{B}_{c}[0.62];h_{b}\bar{B}_{c}[0.57]$
16185.6 1 $\eta_{b}\bar{B}_{c1}[0.65];\Upsilon\bar{B}_{c0}[0.33];\chi_{b1}\bar{B}_{c}[0.39];h_{b}\bar{B}_{c}[0.31]$
Table 6: The quantum numbers, masses, eigenvectors of the basis wave functions and decay channels of $cc][\bar{b}\bar{b}]$ configuration. For certain states, there are no available channels that both satisfy the conservation laws and have thresholds below the mass of the state.
$J^{PC}$ MM(MeV) eigenvector Decay channels $\Gamma$(MeV)]
$0^{+}$ 12552.2 (0.598,0.801) $B_{c}\bar{B}_{c}[0.24]$
12673.9 (0.801,0.598)(-0.801,0.598) $B_{c}\bar{B}_{c}[1.49];B_{c}^{*}\bar{B}_{c}^{*}[0.46]$
$1^{+}$ 12611.6 1 $B_{c}\bar{B}_{c}^{*}[0.33]$
$2^{+}$ 12643.4 1
$0^{-}$ 12983.9 1
$1^{-}$ 12984.6 (0.598,0.801)
13106.3 (0.801,0.598)(-0.801,0.598) $B_{c}\bar{B}_{c1}[1.24];B_{c}^{*}\bar{B}_{c0}[1.01];B_{c}^{*}\bar{B}_{c1}[0.79]$
13013.9 1 $B_{c}\bar{B}_{c1}[0.24]$
12985.6 1
III.2 $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$and $cb][\bar{c}\bar{b}]$ systems
From the wave functions of $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$ and $cb][\bar{c}\bar{b}]$ configurations, it can be seen that these states possesses a well-defined charge conjugation parity ($C$-parity). For these states, conservation of $C$-parity must be taken into account during the decay process. Their spectra and decay channels were shown in Fig. 1 and Tables 2-3. The final-state mesons we choose for these decays are in the $1S$ and $1P$ states. To date, only the spin-0 states of the $B_{c}(1S)$ and $B_{c}(2S)$ mesons have been observed experimentally. Therefore, for the remaining $c\bar{b}$ mesons in the $1S$ and $1P$ states, we use the masses predicted by lattice QCD calculations [58]. For simplicity, we denote the $c\bar{b}$ mesons with quantum numbers $J^{PC}=0^{+}$ and $1^{+}$ as $B_{c0}$ and $B_{c1}$, respectively.
III.2.1 $cc][\bar{c}\bar{c}]$ system
For the tetraquark state $cc][\bar{c}\bar{c}]$ with different quantum numbers, the mass spectra and corresponding meson-meson thresholds are exhibited in Fig. 1(a) and the decay channels are shown in Table 2.
The $S$-wave fully charmed tetraquark states possess quantum numbers $J^{PC}=0^{++}$, $1^{+-}$, and $2^{++}$. There are two states with $0^{++}$, with masses of 6098.45 MeV and 6314.21 MeV, respectively. The lower-mass state, at 6098.45 MeV, lies above both the $\eta_{c}\eta_{c}$ and $\eta_{c}J/\psi$ thresholds. However, due to conservation laws, it can only decay naturally into the $\eta_{c}\eta_{c}$ channel, while decay into $\eta_{c}J/\psi$ is forbidden. The higher-mass $0^{++}$ state, at 6314.21 MeV, lies above both $\eta_{c}\eta_{c}$ and $J/\psi J/\psi$ thresholds and is allowed to decay into both of these two final states. However, the width of the $J/\psi J/\psi$ final state is smaller, indicating that this state is more likely to decay into the $\eta_{c}\eta_{c}$ channel.
For the $1^{+-}$ and $2^{++}$ states, the masses are 6204.49 MeV and 6260.89 MeV, respectively. Both lie above the $\eta_{c}J/\psi$ and $J/\psi J/\psi$ thresholds. Nonetheless, due to quantum number constraints, the $1^{+-}$ state can only decay into the $\eta_{c}J/\psi$ channel, whereas the $2^{++}$ state is allowed to decay solely into the $J/\psi J/\psi$ channel.
The $P$-wave fully charmed tetraquark states have quantum numbers $J^{PC}=0^{-+}$, $1^{--}$, and $1^{-+}$. There is one $0^{-+}$ state, with a mass of 6543.23 MeV, which can decay only into the $\eta_{c}\chi_{c0}$ channel. Three states are identified with $J^{PC}=1^{--}$, having masses of 6524.51 MeV, 6555.97 MeV, and 6740.27 MeV, respectively. The two lower-mass states, at 6524.51 MeV and 6555.97 MeV, lie above the $J/\psi\chi_{c0}$ threshold but below the $J/\psi\chi_{c1}$ threshold. Consequently, due to conservation laws, they are only allowed to decay into the $\eta_{c}h_{c}$ and $J/\psi\chi_{c0}$ channels. The masses of both two states are in good agreement with the mass of the $X(6600)$ observed by the CMS Collaboration, reported as $M=6638^{+43}_{-38}(\text{stat})^{+16}_{-31}(\text{syst})$ MeV [7]. The highest $1^{--}$ state, at 6740.27 MeV, lies above the $J/\psi\chi_{c2}$ threshold and can decay into $\eta_{c}h_{c}$, $J/\psi\chi_{c0}$, $J/\psi\chi_{c1}$, and $J/\psi\chi_{c2}$ final states. This state is more likely to decay into the $\eta_{c}h_{c}$ and $J/\psi\chi_{c0}$ channels due to their larger widths. In addition, there is a single $1^{-+}$ state with a mass of 6586.89 MeV, which can decay only into the $\eta_{c}\chi_{c1}$ channel.
III.2.2 $bb][\bar{b}\bar{b}]$ system
For the tetraquark state $bb][\bar{b}\bar{b}]$ with different quantum numbers, the mass spectra and corresponding meson-meson thresholds are exhibited in Fig. 1(b) and the decay channels are shown in Table 2.
The quantum numbers of the $S$-wave fully bottom tetraquark states are $J^{PC}=0^{++}$, $1^{+-}$, and $2^{++}$. There are two $0^{++}$ states with masses of 18948.9 MeV and 19066.9 MeV. The $1^{+-}$ and $2^{++}$ states have masses of 19006.9 MeV and 19037.7 MeV, respectively. All these $S$-wave states lie above the $\Upsilon\Upsilon$ threshold. Both $0^{++}$ states are allowed to decay into the $\eta_{b}\eta_{b}$ and $\Upsilon\Upsilon$ channels.But this state is more likely to decay into the $\eta_{b}\eta_{b}$ channel due to its larger width. The $1^{+-}$ state can decay only into the $\eta_{b}\Upsilon$ channel, while the $2^{++}$ state is allowed to decay only into the $\Upsilon\Upsilon$ channel.
The quantum numbers of the $P$-wave fully bottom tetraquark states are $J^{PC}=0^{-+}$, $1^{--}$, and $1^{-+}$. Their masses all lie above the $\Upsilon\chi_{b2}$ threshold. Taking conservation laws into account, the $0^{-+}$ state can decay into the $\eta_{b}\chi_{b0}$ and $\Upsilon h_{b}$ channels, while the $1^{-+}$ state is allowed to decay into the $\eta_{b}\chi_{b1}$ and $\Upsilon h_{b}$ channels. The $1^{--}$ state can decay into multiple channels, including $\eta_{b}h_{b}$, $\Upsilon\chi_{b0}$, $\Upsilon\chi_{b1}$, and $\Upsilon\chi_{b2}$.
III.2.3 $cb][\bar{c}\bar{b}]$ system
For the tetraquark state $cb][\bar{c}\bar{b}]$ with different quantum numbers, the mass spectra and corresponding meson-meson thresholds are exhibited in Fig. 1(c) and the decay channels are shown in Table 3.
Since the $cb][\bar{c}\bar{b}]$ tetraquark system contains no identical quarks or antiquarks, it is not subject to the Pauli exclusion principle. As a result, there is no restriction on the construction of basis states, and all possible spin-parity and charge-conjugation quantum numbers can be realized. This leads to a rich spectra of possible states, including $0^{++}$, $^{++}, 1+1^{+-}$, $2^{++}$ for $S$-wave and $^{--}, 0+0^{-+}$, $^{--}, 1+1^{-+}$ for $P$-wave.
There are four states with quantum numbers $0^{++}$, with masses of 12477.8 MeV, 12548.6 MeV, 12633.4 MeV, and 12690.1 MeV, respectively. Among them, the two lighter states, 12477.8 MeV and 12548.6 MeV, lie above the $\eta_{c}\eta_{b}$ threshold but below the $J/\psi\Upsilon$ threshold. Therefore, considering the conservation of angular momentum and $C$-parity, they can only decay into the $\eta_{c}\eta_{b}$ channel. In addition, the $cb][\bar{c}\bar{b}]$ tetraquark states can decay not only into $c\bar{c}$ and $b\bar{b}$ mesons, but also into $c\bar{b}$ and $b\bar{c}$ mesons. However, since the masses of the two lighter states are below the $B_{c}\bar{B}_{c}$ threshold, decays into $c\bar{b}$ and $b\bar{c}$ mesons are kinematically forbidden. The two heavier states, with masses of 12633.4 MeV and 12690.1 MeV, lie above the $J/\psi\Upsilon$ threshold and can thus decay into both $\eta_{c}\eta_{b}$ and $J/\psi\Upsilon$ channels. Moreover, since their masses are also above the $B_{c}\bar{B}_{c}$ threshold, decays into this channel are allowed. In particular, the state with mass 12690.1 MeV lies above the $B_{c}^{*}\bar{B}_{c}^{*}$ threshold, and therefore can also decay into the $B_{c}^{*}\bar{B}_{c}^{*}$ channel. The two states, 12633.4 MeV and 12690.1 MeV, are more likely to decay into $\eta_{c}\eta_{b}$ channel due to its larger width.
There are two states with quantum numbers $1^{++}$, with masses of 12582.8 MeV and 12665.6 MeV, respectively. Both states lie above the $J/\psi\Upsilon$ threshold and are therefore allowed to decay into this channel. In addition, the heavier state, with a mass of 12665.6 MeV, lies above the thresholds for both $B_{c}\bar{B}_{c}^{*}$ and $B_{c}^{*}\bar{B}_{c}^{*}$, making decays into these channels kinematically allowed as well. But the width of $J/\psi\Upsilon$ channel is larger than $B_{c}\bar{B}_{c}^{*}$ and $B_{c}^{*}\bar{B}_{c}^{*}$, thus the state of 12665.6 MeV is more likely to decay into $J/\psi\Upsilon$ channel.
There are four states with quantum numbers $1^{+-}$, with masses of 12547.4 MeV, 12590.8 MeV, 12617.6 MeV, and 12655.4 MeV, respectively. All of these states lie above the $\eta_{c}\Upsilon$ and $J/\psi\eta_{b}$ thresholds, allowing them to decay into both channels. Additionally, the two heavier states, with masses of 12617.6 MeV and 12655.4 MeV, are above the $B_{c}\bar{B}_{c}^{*}$ threshold and can therefore also decay into this channel. Both the four $1^{+-}$ states are more likely to decay into $\eta_{c}\Upsilon$ channel because the width of $\eta_{c}\Upsilon$ is larger than that of other channels.
There are two states with quantum numbers $2^{++}$, whose masses are 12649.1 and 12666.1 MeV, respectively. Both them lie above the $J/\psi\Upsilon$ threshold and are allowed to decay into this channel. Furthermore, the state with 12666.1 MeV lies above the threshold of $B_{c}^{*}\bar{B}_{c}^{*}$, therefore can also decay into the $B_{c}^{*}\bar{B}_{c}^{*}$ channel. The state of 12666.1 MeV is more likely to decay into $J/\psi\Upsilon$ channel due to the larger width.
There are two $0^{--}$ states, whose masses are 12955.1 MeV and 13037.9 MeV. The mass of 13037.9 MeV lies above $J/\psi\chi_{b1}$, $\chi_{c1}\Upsilon$ and $B_{c}\bar{B}_{c0}$ thresholds therefore this state can naturally decay into these channels. The state with a mass of 12955.1 MeV has no available $S$-wave decay channels. Because its mass lies below $J/\psi\chi_{b1}$, $\chi_{c1}\Upsilon$ and $B_{c}\bar{B}_{c0}$ thresholds. However, it can decay into the $J/\psi\Upsilon$ channel via a $P$-wave transition.
There are four $0^{-+}$ states with masses of 12919.7 MeV, 12963.1 MeV, 12989.9 MeV and 13027.7 MeV. Their masses lie all above the thresholds of $\eta_{c}\chi_{b0}$ and $\chi_{c0}\eta_{b}$, thus they can naturally decay into these two channels. The heaviest state lies above the thresholds of the $J/\psi h_{b}$, $h_{c}\Upsilon$, and $B_{c}\bar{B}_{c0}$ channels, and can therefore decay spontaneously into all three of them. Both the four states are more likely to decay into $\chi_{c0}\eta_{b}$ channel due to the larger width.
There are eight $P$-wave tetraquark states with $J^{PC}=1^{--}$. Their masses and decay channels are exhibited in Table 3. It is worth noting that the state with a mass of 13067.9 MeV lies very close to the $B_{c}^{*}\bar{B}_{c1}$ threshold. As a result, it can easily decay into this channel and is expected to appear as a narrow resonance due to the proximity to the threshold.
There are four $1^{-+}$ states with masses of 12949.7 MeV, 12993.1 MeV, 13019.9 MeV, and 13057.7 MeV. All of these states lie above the $\eta_{c}\chi_{b1}$ and $\chi_{c1}\eta_{b}$ thresholds, and can therefore naturally decay into these two channels. In addition, the state with a mass of 12993.1 MeV also lies above the $h_{c}\Upsilon$ threshold, making this decay channel accessible as well. The two higher-mass states, 13019.9 MeV and 13057.7 MeV, lie above even higher thresholds, including $\chi_{c1}\eta_{b}$ and $B_{c}\bar{B}_{c1}$. Furthermore, the heaviest state, with a mass of 13057.7 MeV, lies above the $B_{c}^{*}\bar{B}_{c0}$ threshold and can decay into this channel as well.
III.3 $cc][\bar{c}\bar{b}]$, $bb][\bar{b}\bar{c}]$ and $cc][\bar{b}\bar{b}]$ systems
For the $cc][\bar{c}\bar{b}]$, $bb][\bar{b}\bar{c}]$, and $cc][\bar{b}\bar{b}]$ systems, $C$-parity is not well-defined due to their flavor asymmetry. As a result, the possible quantum numbers for these configurations are $J^{P}=0^{+}$, $1^{+}$, $2^{+}$, $0^{-}$, and $1^{-}$. Our theoretical results are shown in Fig. 2 and Tables 4-6. The analysis method is similar to that for the $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$and $cb][\bar{c}\bar{b}]$ tetraquark states.
IV Summary
In this paper, we systematically study the mass spectra of the lower $S$- and $P$-wave fully heavy tetraquark states, as well as analysis their possible strong decay channels. Considering the basic features of the fully heavy tetraquark system, we adopted an improved Hamiltonian based on the one gluon exchange process. The Hamiltonian include constituent quark masses, spin-spin interaction, spin-orbit coupling and pure orbital contribution. The model parameters required for our calculations are derived from fitting conventional mesons and baryons. The parameters corresponding to different color structures can be calculated through $SU(3)$ group algebra.
For the fully charmed tetraquark system, our results do not support interpreting $X(6600)$, $X(6900)$, and $X(7200)$ as $S$-wave fully charmed tetraquark states. It is worth noting that we predict two fully charmed tetraquark states with quantum numbers $J^{PC}=1^{--}$, whose masses are 6524.51 MeV and 6555.97 MeV respectively. At the same time, we analyzed the decay states of these two states and found that their characteristics are consistent with the observations of the CMS Collaboration on the $X(6600)$ resonance. For the fully bottom tetraquark system, our calculated masses are all above the $\Upsilon\Upsilon$ threshold and exceed the $bb\bar{b}\bar{b}$ mass reported by CMS Collaboration and ANDY Collaboration. Nevertheless, LHCb Collaboration and other experiments have not observed any $bb\bar{b}\bar{b}$ resonant states thus far. Further experimental investigations are required to address this question. In addition, we calculated masses of several other fully heavy tetraquark states and analysis their strong decay channels. It is also anticipated that some of the states predicted in this work may be observed in future experiments.
Acknowledgements.
The authors would like to thank Yanmei Xiao for helpful comments and discussions. This work was supported by the Scientific Research Foundation of Chengdu University of Technology under grant No.10912-KYQD2022-09557.
Appendix A Wave Functions
The wave functions of all fully heavy tetraquark states are given below.
[…]
$\begin{gathered}|0^{+(+)}\rangle_{1}=|(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{\bar{6}}\rangle_{0},\\ |0^{+(+)}\rangle_{2}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{0},\end{gathered}$ (16)
$|1^{+(-)}\rangle=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{1},$ (17)
$|2^{+(+)}\rangle=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{2},$ (18)
$|0^{-(+)}\rangle=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{1},$ (19)
$\begin{gathered}|{1^{-(-)}}\rangle_{1}=|{(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{\bar{6}}}\rangle_{0},\\ |{1^{-(-)}}\rangle_{2}=|{(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}}\rangle_{0},\\ |{1^{-(+)}}\rangle_{3}=|{(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}}\rangle_{1},\\ |{1^{-(-)}}\rangle_{4}=|{(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}}\rangle_{2}.\end{gathered}$ (20)
[…]
$\begin{gathered}|{0^{+}}\rangle_{1}=|{(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{\bar{6}}}\rangle_{0},\\ |{0^{+}}\rangle_{2}=|{(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}}\rangle_{0},\end{gathered}$ (21)
$\begin{gathered}|1^{+}\rangle_{1}=|(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{\bar{6}}\rangle_{1},\\ |1^{+}\rangle_{2}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{3}\rangle_{1},\\ |1^{+}\rangle_{3}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{1},\end{gathered}$ (22)
$|2^{+}\rangle=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{2}$ (23)
$\begin{gathered}|0^{-}\rangle_{1}=|(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{\bar{6}}\rangle_{1},\\ |0^{-}\rangle_{2}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{3}\rangle_{1},\\ |0^{-}\rangle_{3}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{1},\end{gathered}$ (24)
$\begin{gathered}|1^{-}\rangle_{1}=|(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{\bar{6}}\rangle_{0},\\ |1^{-}\rangle_{2}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{0},\\ |1^{-}\rangle_{3}=|(Q_{1}Q_{2})_{0}^{6}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{\bar{6}}\rangle_{1},\\ |1^{-}\rangle_{4}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{0}^{3}\rangle_{1},\\ |1^{-}\rangle_{5}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{1},\\ |1^{-}\rangle_{6}=|(Q_{1}Q_{2})_{1}^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})_{1}^{3}\rangle_{2}.\end{gathered}$ (25)
(3) $cb\bar{c}\bar{b}$
$\begin{gathered}|0^{++}\rangle_{1}=|(cb)_{0}^{6}(\bar{c}\bar{b})_{0}^{\bar{6}}\rangle_{0},\\ |0^{++}\rangle_{2}=|(cb)_{0}^{\bar{3}}(\bar{c}\bar{b})_{0}^{3}\rangle_{0},\\ |0^{++}\rangle_{3}=|(cb)_{1}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{0},\\ |0^{++}\rangle_{4}=|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{0},\end{gathered}$ (26)
$\begin{gathered}|1^{++}\rangle_{1}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{6}(\bar{c}\bar{b})_{0}^{\bar{6}}\rangle_{1}+|(cb)_{0}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1}),\\ |1^{++}\rangle_{2}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{0}^{3}\rangle_{1}+|(cb)_{0}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1}),\end{gathered}$ (27)
$\begin{gathered}|1^{+-}\rangle_{1}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{6}(\bar{c}\bar{b})_{0}^{\bar{6}}\rangle_{1}-|(cb)_{0}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1}),\\ |1^{+-}\rangle_{2}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{0}^{3}\rangle_{1}-|(cb)_{0}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1}),\\ |1^{+-}\rangle_{3}=|(cb)_{1}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1},\\ |1^{+-}\rangle_{4}=|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1},\end{gathered}$ (28)
$\begin{gathered}|2^{++}\rangle_{1}=|(cb)_{1}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{2},\\ |2^{++}\rangle_{2}=|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{2},\end{gathered}$ (29)
$\begin{gathered}|0^{--}\rangle_{1}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{6}(\bar{c}\bar{b})_{0}^{\bar{6}}\rangle_{1}+|(cb)_{0}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1}),\\ |0^{--}\rangle_{2}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{0}^{3}\rangle_{1}+|(cb)_{0}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1}),\end{gathered}$ (30)
$\begin{gathered}|0^{-+}\rangle_{1}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{6}(\bar{c}\bar{b})_{0}^{\bar{6}}\rangle_{1}-|(cb)_{0}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1}),\\ |0^{-+}\rangle_{2}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{0}^{3}\rangle_{1}-|(cb)_{0}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1}),\\ |0^{-+}\rangle_{3}=|(cb)_{1}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1},\\ |0^{-+}\rangle_{4}=|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1},\end{gathered}$ (31)
$\begin{gathered}|1^{--}\rangle_{1}=|(cb)_{0}^{6}(\bar{c}\bar{b})_{0}^{\bar{6}}\rangle_{0},\\ |1^{--}\rangle_{2}=|(cb)_{0}^{\bar{3}}(\bar{c}\bar{b})_{0}^{3}\rangle_{0},\\ |1^{--}\rangle_{3}=|(cb)_{1}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{0},\\ |1^{--}\rangle_{4}=|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}v_{0},\\ |1^{--}\rangle_{5}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{6}(\bar{c}\bar{b})_{0}^{\bar{6}}\rangle_{1}+|(cb)_{0}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1}),\\ |1^{--}\rangle_{6}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{0}^{3}\rangle_{1}+|(cb)_{0}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1}),\\ |1^{--}\rangle_{7}=|(cb)_{1}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{2},\\ |1^{--}\rangle_{8}=|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{2},\end{gathered}$ (32)
$\begin{gathered}|1^{-+}\rangle_{1}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{6}(\bar{c}\bar{b})_{0}^{\bar{6}}\rangle_{1}-|(cb)_{0}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1}),\\ |1^{-+}\rangle_{2}=\frac{1}{\sqrt{2}}(|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{0}^{3}\rangle_{1}-|(cb)_{0}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1}),\\ |1^{-+}\rangle_{3}=|(cb)_{1}^{6}(\bar{c}\bar{b})_{1}^{\bar{6}}\rangle_{1},\\ |1^{-+}\rangle_{4}=|(cb)_{1}^{\bar{3}}(\bar{c}\bar{b})_{1}^{3}\rangle_{1},\end{gathered}$ (33)